# A High-Performance Evolutionary Multiobjective Community Detection Algorithm

### v1

@@ 3 The HP-MOCD Algorithm
This section details the High-Performance Multi-Objective Community Detection (HP-MOCD) algorithm, a scalable evolutionary method for efficiently identifying high-quality community partitions in large complex networks. HP-MOCD combines the NSGA-II optimization framework with a parallel architecture and topology-aware genetic operators tailored to the structure of real-world graphs. In addition to detailing its core components, we describe the algorithm’s design choices, solution representation, and multi-objective selection strategy. The implementation is written in Rust for performance and exposed to Python via PyO3. The full source code is publicly available on GitHub.
@@ 3.1 Overview and Design Rationale
Let us consider a network represented by $G=(V,E)$. The goal of HP-MOCD is to uncover community structures in $G$ by minimizing the two objectives proposed by [7] in Equation 6. By jointly optimizing these criteria, the algorithm promotes partitions in which communities are internally cohesive and externally well-separated, balancing internal density with structural isolation. To accomplish this, the algorithm builds upon the NSGA-II framework, a widely adopted and effective method in the field of multi-objective evolutionary algorithms. NSGA-II was chosen primarily due to its strong performance in producing well-distributed and convergent Pareto fronts, especially when compared to earlier algorithms like PESA-II [29], which tend to suffer from loss of diversity or weaker selection pressure.
The HP-MOCD procedure unfolds in two main phases. First, it constructs an initial population of potential solutions (called individuals), each representing a possible way to partition the graph into communities. Then, it evolves this population over a number of generations using genetic operators, continuously selecting and refining the best trade-offs between the objectives. This process is summarized in Algorithm 1.
Algorithm 1 Overview of the HP-MOCD Algorithm.
1:procedure HP-MOCD(Graph $G$, Population Size $N$, Max Generations $T$, Crossover Probability $C_{P}$, Mutation Probability $M_{P}$)
2:     $P\leftarrow\text{InitializePopulation}(G,N)$
@@
4:     for $gen\leftarrow 0$ to $T$ do
5:         ComputeCrowdingDistance$(P)$
6:         $M\leftarrow\text{TournamentSelection}(P,N)$triangleright By rank & crowding
7:         $Q\leftarrow\text{CreateOffSpring}(M,\ G,\ C_{P},\ M_{P},4)$triangleright (Algorithm 2)
8:         EvaluatePopulation$(Q,\ G)$
9:         $R\leftarrow P\cup Q$
10:         $P\leftarrow\text{SelectNextGeneration}(R,N)$triangleright Keep best $N$ by rank + crowding
11:     end for
@@
12:     $\mathcal{F}_{1}\leftarrow\{\,x\in P\mid\text{rank}(x)=1\}$
13:     return $\mathcal{F}_{1}$
14:end procedure
Internally, the graph $G$ is stored using a high-performance hash map that maps each node to its list of neighbors. This representation, based on a fast, non-cryptographic hashing scheme  [^2] A custom hash algorithm was used by rustc (plus hashmap/set aliases): fast, deterministic, not secure https://github.com/rust-lang/rustc-hash, enables efficient access and mutation operations during the evolution cycle, which is especially important for large-scale networks. In the subsequent phase, the NSGA-II algorithm produces a set of non-dominated solutions that collectively form a Pareto front. A Pareto front comprises multiple optimal solutions, each offering a different trade-off between the two conflicting objectives. A solution, commonly referred to as an individual in evolutionary algorithms, is specifically represented as a mapping (hashmap) that associates each node of the graph with a particular community identifier. Thus, each individual represents a unique community partitioning of the graph.
To better illustrate this concept, consider the toy network shown in Figure 1. Based on this graph, HP-MOCD may generate many different community configurations over time. These solutions reflect different compromises between internal density and external sparsity, and none is necessarily superior to the others in all aspects.
@@
Figure 1: A simple network with some nodes that shows aspects of a community.
All these individuals belong to the final Pareto front produced by the algorithm. Since no single partition can simultaneously minimize both objectives without compromising the other, the algorithm instead returns a diverse set of non-dominated solutions. Each one represents a valid alternative, offering different structural insights. An individual (solution) produced by our algorithm is stored as a mapping,
$\mathcal{M}:V\;\longrightarrow\;\{A,B,C,\dots\},$
For example, consider the decoded individual for the graph in Figure 1:
@@
If we choose $\phi(A)=\text{blue}$, $\phi(B)=\text{green}$, and $\phi(C)=\text{red}$, then nodes $\{1,2,3,4\}$ appear in blue, $\{5,6,7,8\}$ in green, and $\{9,10,11,12\}$ in red, as shown in Figure 2. And this can be just one of a large set of solutions that the proposed HP-MOCD can return to us.
123456789101112
Figure 2: The same network, but with distinct communities. Nodes within each community (blue, green, and red) are more densely connected internally, with sparser connections between communities.
This example illustrates a key advantage of the multi-objective approach: instead of converging to a single solution, HP-MOCD yields a diverse set of high-quality alternatives. This flexibility allows to select the partition that best fits the specific needs or constraints of their application domain.
@@ 3.2.1 Crossover
The crossover operator is responsible for combining multiple parent solutions to generate a new offspring. Let $P=(P_{1},\ldots,P_{n})$ denote a collection of $n$ parent individuals, where each individual $P_{i}$ represents a partition of the graph $G=(V,E)$, that is, a mapping from each node $v\in V$ to a community label $c\in C$. The operator is controlled by a crossover threshold parameter $C_{R}\in[0,1]$. Initially, a random number $x$ is sampled uniformly from the interval $[0,1]$. If $x>C_{R}$, the crossover is skipped, and one of the parent partitions is selected uniformly at random to serve as the offspring:
$x>C_{R}\quad\Rightarrow\quad P_{\text{child}}=\text{Random}(P_{1},\ldots,P_{n}).$ (7)
@@
$\text{count}(c,v)=\sum_{i=1}^{n}\mathbf{1}\{P_{i}(v)=c\},$ (8)
where $\mathbf{1}\{\cdot\}$ denotes the indicator function. The community (or communities) with the highest count is then identified as
@@
In the event of a tie, one of the top-scoring communities is selected uniformly at random. The selected label $c^{*}(v)$ is then assigned to node $v$ in the offspring partition, that is, $P_{\text{child}}(v)=c^{*}(v)$. This procedure is applied to all nodes in $V$, producing a complete offspring partition. The resulting solution tends to preserve dominant structural features from the parents while still allowing for variability and exploration through random tie-breaking.
@@ 3.2.2 Mutation
The community assignment of $v$ is updated to this most common community. After all selected nodes have been processed, the original mapping is updated with these new community assignments. The Offspring Generation procedure, which is presented in detail in Algorithm 2, is performed as follows.
Algorithm 2 Offspring Generation
1:procedure CreateOffspring($M$, $G$, $C_{P}$, $M_{P}$, $E_{S}$)
2:     $Q\leftarrow[\ ]$triangleright Initialize offspring population
3:     for $i\leftarrow 1$ to $|M|$ do
4:         $x\leftarrow\text{random}(0,1)$
5:         if $x<C_{P}$ then $\triangleright$ Use $E_{s}$ parents to crossover
6:              $Parents\leftarrow\text{SelectRandom}(M,E_{S})$
7:              $P_{child}\leftarrow\text{Crossover}(Parents)$
8:         else
9:              $P_{child}\leftarrow\text{Random}(M)$
10:         end if
11:         Mutate$(P_{child},G,M_{P})$
12:         Append $P_{child}\rightarrow Q$
13:     end for
14:     return $Q$
15:end procedure
3.2.3 Parameters
The evolutionary process in HP-MOCD is governed by two key hyperparameters: population size and number of generations. To determine appropriate values, we conducted empirical experiments on synthetic benchmarks, using the same configuration as in our main experiments, except for varying $n$ (see Section 4.1.1). We analyzed the convergence behavior of the algorithm under these settings. As shown in Figure 3, 100 generations provided a favorable trade-off between solution quality and computational cost; for the other parameters and scenarios, we will leave them in the supplementary material (See References).
(a) 10 Generations
(b) 30 Generations
(c) 50 Generations
(d) 80 Generations
(e) 100 Generations
(f) 110 Generations
Figure 3: Performance results across varying $maxgen$ values with $\mu=0.1$. Refer to Section 4.1.1 for details on the experimental setup.
3.3 Pareto Selection
As discussed in Section 2, multi-objective optimization algorithms such as NSGA-II do not return a single “optimal” solution. Instead, they produce a Pareto front, a set of non-dominated solutions, in which no solution is strictly better than another across all objectives. This naturally raises the question: how can one select the most appropriate solution from the Pareto front?
@@ 3.3 Pareto Selection
$Q(C)=1-\text{intra}(C)-\text{inter}(C)$ (11)
This score favors solutions with low intra-community penalty and low inter-community connectivity, as defined in Equations 6, thus promoting partitions that are both cohesive and well-separated. Figure 4 illustrates this behavior: the red dot indicates the solution that maximizes $Q(C)$. We observe that this selection strategy consistently identifies partitions that also perform well under widely used external validation metrics such as modularity, NMI, and AMI.
@@ 3.4 Complexity Analysis
Complexity time of HP-MOCD per generation is driven by its genetic operators (crossover and mutation) and the key NSGA-II routines (non-dominated sorting and crowding distance). Let $|V|$ be the number of nodes, $|E|$ the number of edges, $N_{p}$ the population size, and $M_{R}$ the mutation rate.
@@ 3.4.1 Operators Complexity Analysis
For each node in the graph, the algorithm crossover identifies the most frequent community label assigned to that node among a fixed number of parent solutions. Since hash map operations have an expected constant time complexity, $O(1)$, processing each node takes $O(1)$ time. Assigning the most frequent label also takes constant time per node. Consequently, for all $|V|$ nodes, the expected time complexity of a single crossover operation is linear in the number of nodes, i.e., $O(|V|)$.
The mutation iterates through each of the $|V|$ nodes, selecting a node for mutation with a given probability $M_{R}\in[0,1]$. This selection phase involves a random draw for each node, costing $O(|V|)$ time, and, for each node selected for mutation, inspects its local neighborhood to determine a new community assignment (by finding the most frequent community label among its neighbors). This requires visiting all adjacent nodes, so the cost for mutating a single node $v$ is proportional to its degree, $d(v)$. The expected number of nodes selected for mutation is $M_{R}\cdot|V|$. The total time for recalculating assignments for these mutated nodes is proportional to the sum of their degrees. In expectation, this sum is $M_{R}\sum_{v\in V}d(v)=M_{R}\cdot 2|E|$. Therefore, the expected time complexity for this part of the mutation is $O(M_{R}\cdot|E|)$. Combining the node selection phase and the assignment update phase, the total time complexity for one mutation operation is $O(|V|+M_{R}\cdot|E|)$. For sparse graphs, where $|E|=O(|V|)$, and assuming $M_{R}$ is a constant, this simplifies to $O(|V|)$.
@@
The dominant cost per generation is the sum of: $T_{\text{gen}}=O(N_{p}|V|)+O(N_{p}\log N_{p})+O(N_{p}\log N_{p})$ which simplifies to $T_{\text{gen}}=O(N_{p}|V|+N_{p}\log N_{p})$. Since $|V|$ in $G$ is typically much larger than $\log N_{p}$, the $O(N_{p}|V|)$ term usually dominates. Thus, the complexity per generation effectively becomes: $O(N_{p}|V|)$ for sparse graphs. Over $G$ generations, the total time complexity of HP-MOCD is: $O(G\cdot N_{p}|V|)$.
4 Experiments
This section presents a series of experiments designed to assess the performance of the proposed HP-MOCD algorithm in large-scale networks. Our goal is to evaluate both the quality of the detected communities and the scalability of the method in terms of runtime. All experiments are fully reproducible, with code and data generation scripts publicly available on GitHub.
4.1 Evaluation Setup
To evaluate the performance of the proposed HP-MOCD algorithm, we conducted a set of experiments using synthetic benchmark networks. This setup enables the controlled variation of structural properties, such as size, noise, and community separability, while providing known ground-truth partitions for external validation.
4.1.1 Synthetic Benchmark Networks
We used the LFR benchmark generator [30], implemented via the NetworkX Python library, to generate realistic artificial networks with built-in community structure. These benchmarks are widely used due to their ability to replicate real-world network features, such as power-law degree distributions and heterogeneous community sizes. Two main experimental settings were considered:
  • Scalability test: The number of nodes $n$ varied from 10,000 to 100,000, while all other parameters remained fixed. This setup evaluates how each algorithm scales computationally and structurally with increasing network size.
  • Robustness test: With network size fixed at $n=100{,}000$, the mixing parameter $\mu\in[0.1,0.8]$ was varied. This parameter controls the proportion of each node’s links that connect to other communities. Smaller values of $\mu$ correspond to well-separated communities, while higher values represent noisy, overlapping structures that are harder to detect.
All LFR graphs were generated using the following fixed parameters: maximum degree = 50, average degree = 20, minimum community size = 20, maximum community size = 100, degree exponent $\tau_{1}=2.5$, and community size exponent $\tau_{2}=1.5$. It is important to note that in all generated synthetic networks, the proportion of overlapping nodes and the number of communities per overlapping node were both set to zero. This choice reflects our focus on non-overlapping community detection. The known ground truth structure of these LFR benchmark networks allowed us to directly measure and evaluate the accuracy and effectiveness of the HP-MOCD algorithm.
4.1.2 Algorithms and Parameter Configuration

### v2

@@ 4 The HP-MOCD Algorithm
We now describe the High-Performance Multi-Objective Community Detection (HP-MOCD) algorithm. Our method integrates the NSGA-II optimization framework with a parallel architecture and custom genetic operators designed for network topology. The implementation is written in Rust for performance and exposed to Python via PyO3. The full source code is publicly available on GitHub and the documention and usage are available on the Official Documentation. Algorithm 1 and Figure 2 summarizes the HP-MOCD’s algorithm.
@@ 4.1 Overview and Design Rationale
Traditional single-objective approaches often suffer from resolution limit problems and can become trapped in local optima, as discussed in Section 2. To address these limitations, we need to use metrics that discover the entire Pareto front of solutions, ranging from coarse to fine partitions without biasing toward any particular resolution level.
Given a graph $G=(V,E)$ with node set $V$ and edge set $E$, let $C=\{c_{1},c_{2},\ldots,c_{k}\}$ represent a partition of $V$ into $k$ communities. The set of all valid partitions of $V$ is denoted as the feasible space $\Omega=\{C\mid C\text{ is a valid partition of }V\}$.
Following the formulation of [7], we decompose the community detection problem into two complementary objectives that capture the essential trade-offs in community structure. Our method seeks to find partitions that simultaneously minimize both objective functions:
$f_{1}(C)=1-\sum_{c\in C}\frac{\lvert E(c)\rvert}{m},\quad f_{2}(C)=\sum_{c\in C}\Bigl{(}\frac{\sum_{v\in c}\deg(v)}{2m}\Bigr{)}^{2}$ (6)
@@
The first objective function $f_{1}$ (or $intra(C)$) measures the fraction of edges that are not internal to communities, effectively penalizing partitions with poor internal cohesion. Minimizing $f_{1}$ encourages the formation of densely connected communities. The second objective function $f_{2}$ (or $inter(C)$) represents the sum of squared relative community sizes (measured by total degree), which penalizes the formation of overly large communities. Minimizing $f_{2}$ promotes more balanced partitions and prevents the emergence of dominant communities that absorb an excessive number of nodes.
These objectives are inherently conflicting; optimizing for strong internal cohesion (low $f_{1}$) often leads to fewer, larger communities, while optimizing for balanced community sizes (low $f_{2}$) typically results in many smaller communities with potentially weaker internal structure.
The problem is then stated as:
$\min_{C\,\in\,\Omega}\;\mathbf{f}(C)\;=\;\min_{C\,\in\,\Omega}\;\bigl{(}f_{1}(C),\,f_{2}(C)\bigr{)}.$ (7)
Figure 2 illustrates the diagram of the proposed HP-MOCD algorithm, the initialization, evaluation, and evolutionary selection phases. Parallelism is applied during the evaluation of objectives and sorting operations, enhancing scalability. The evolutionary loop iteratively generates offspring, combines populations, and applies environmental selection until the maximum number of generations is reached.
[…]
Figure 2: Flowchart of the HP-MOCD algorithm. Green-highlighted boxes indicate parallel operations that enhance computational efficiency. The algorithm iteratively evolves populations through selection, crossover, mutation, and environmental selection until convergence.
Our HP-MOCD algorithm addresses this optimization problem, employing the NSGA-II framework and parallel architecture. The overall process is outlined in Algorithm 1. It begins by creating an initial population of random partitions, each representing a possible way to partition the graph into communities. Then, through successive generations, it employs genetic operators (selection, crossover, and mutation) to evolve the population, preserving the best non-dominated solutions.
Algorithm 1 High-Performance Multi-Objective Community Detection (HP-MOCD)
1:procedure HP-MOCD(Graph $G(V,E)$, Population Size $N_{p}$, Max Generations $T$, Crossover Probability $C_{P}$, Mutation Probability $M_{P}$, Ensemble Size $E_{S}$)
[…]
3:  Output: Pareto front $\mathcal{F}_{1}$ of non-dominated community partitions
4:  $\triangleright$ Phase 1: Initialization and Preprocessing
5:  $degrees\leftarrow\text{PrecomputeDegrees}(G)$triangleright $O(|V|)$ cache for $\deg(v),\forall v\in V$
6:  $m\leftarrow|E|$triangleright Total number of edges
7:  $P^{(0)}\leftarrow\text{InitializePopulation}(G,N_{p})$triangleright Random hash-map partitions
8:  $\triangleright$ Phase 2: Initial Population Evaluation
9:  parallel for each $p\in P^{(0)}$ do
[…]
11:  end parallel
12:  $\triangleright$ Phase 3: Evolutionary Optimization Loop
13:  for $gen\leftarrow 1$ to $T$ do
14:   $\triangleright$ Multi-objective selection and ranking
[…]
18:   $\triangleright$ Parallel offspring generation
19:   parallel map $i\leftarrow 1$ to $N_{p}$ do
[…]
23:    return child
24:   end parallel $\rightarrow Q^{(gen)}$
25:   $\triangleright$ Parallel offspring evaluation
26:   parallel for each $q\in Q^{(gen)}$ do
[…]
28:   end parallel
29:   $\triangleright$ Environmental selection
30:   $R^{(gen)}\leftarrow P^{(gen-1)}\cup Q^{(gen)}$triangleright Combine parent and offspring populations
31:   $P^{(gen)}\leftarrow\text{EnvironmentalSelection}(R^{(gen)},N_{p})$triangleright Select best $N_{p}$ individuals
32:  end for
33:  $\triangleright$ Phase 4: Pareto Front Extraction and Solution Selection
@@
34:  $\mathcal{F}_{1}\leftarrow\{p\in P^{(T)}\mid\text{rank}(p)=1\}$triangleright Extract Pareto front
35:  $p^{*}\leftarrow\text{SelectSolution}(\mathcal{F}_{1})$triangleright Algorithm 3
36:  return $\mathcal{F}_{1},p^{*}$triangleright Return Pareto front and recommended solution
37:end procedure
4.2 Solution Representation
Consider the toy network $G(V,E)$ shown in Figure 3(a). The graph representation converts a NetworkX or igraph network into an optimized data structure that maintains multiple complementary representations of the same graph data: (i) An edge list for sequential access. (ii) An adjacency list for neighborhood queries. (iii) Pre-computed degree mappings. (iv) An edge lookup hash set that enables constant-time edge existence queries. This design allows significant performance gains, providing $O(1)$ access times for operations.
@@ 4.2 Solution Representation
123456789101112
Figure 3: (a) A simple network with some nodes that shows aspects of a community. (b) The same network, but with distinct communities. Nodes within each community (blue, green, and red) are more densely connected internally, with sparser connections between communities.
An individual in the population is a candidate solution, represented as a hash map that maps each node $v\in V$ to a community identifier, as a mapping $\mathcal{M}$.
$\mathcal{M}:V\;\longrightarrow\;\{c_{1},c_{2}\dots\,c_{k}\}$
@@
This representation is memory-efficient, allowing for fast lookups. Our custom crossover and mutation operators drive the evolution of these individuals.
@@ 4.3.1 Crossover
The crossover operator is responsible for combining multiple parent solutions to generate a new offspring. Let $P=(P_{1},\ldots,P_{N_{p}})$ denote a collection of ${N_{p}}$ parent individuals, where each individual $P_{i}$ represents a partition of the graph $G=(V,E)$, that is, a mapping from each node $v\in V$ to a community label $c\in C$. The operator is controlled by a crossover threshold parameter $C_{R}\in[0,1]$. Initially, a random number $x$ is sampled uniformly from the interval [0,1][0,1][ 0 , 1 ]. If $x>C_{R}$, the crossover is skipped, and one of the parent partitions is selected uniformly at random to serve as the offspring:
$x>C_{R}\quad\Rightarrow\quad P_{\text{child}}=\text{Random}(P_{1},\ldots,P_{N_{p}}).$ (8)
@@
$\text{count}(c,v)=\sum_{i=1}^{{N_{p}}}\mathds{1}\{P_{i}(v)=c\},$ (9)
where ${}\mathds{1}\{\cdot\}$ denotes the indicator function. The community (or communities) with the highest count is then identified as
@@
In the event of a tie, one of the top-scoring communities is selected uniformly at random. The selected label $c^{*}(v)$ is then assigned to node $v$ in the offspring partition, that is, $P_{\text{child}}(v)=c^{*}(v)$. This procedure is applied to all nodes in $V$, producing a complete offspring partition. The resulting solution tends to preserve dominant structural features from the parents while still allowing for variability and exploration through random tie-breaking. The crossover process is invoked for each of the ${N_{p}}$ offspring to be created in the new population, as detailed in the loop of Algorithm 1 (lines 19-24), thus ensuring diversity and continuous exploration of the search space.
Algorithm 2 Crossover Operator
[…]
2:  Input: Collection of $N_{p}$ parent individuals $P=(P_{1},\ldots,P_{N_{p}})$, crossover threshold $C_{R}\in[0,1]$
3:  Output: Offspring partition $P_{\text{child}}$
4:  Sample $x\sim\text{Uniform}[0,1]$
5:  if $x>C_{R}$ then
6:   return $\text{Random}(P_{1},\ldots,P_{N_{p}})$triangleright Skip crossover, return random parent
7:  end if
8:  Initialize $P_{\text{child}}\leftarrow\emptyset$triangleright Empty offspring mapping
9:  $\triangleright$ Aggregate community assignments from all parents
10:  for each node $v\in V$ do
11:   $\triangleright$ Count occurrences of each community label for node $v$
12:   for each community $c\in C$ do
13:     $\text{count}(c,v)\leftarrow\sum_{i=1}^{N_{p}}\mathds{1}\{P_{i}(v)=c\}$
14:   end for
15:   $\triangleright$ Find community(ies) with highest count
16:   $c^{*}(v)\leftarrow\arg\max_{c\in C}\text{count}(c,v)$triangleright Random tie-breaking if multiple maximizers
17:   $P_{\text{child}}(v)\leftarrow c^{*}(v)$triangleright Assign selected community to node $v$
18:  end for
19:  return $P_{\text{child}}$
20:end procedure
@@ 4.3.2 Mutation
4.4 Solution Selection
For the multi-objective optimization problem formulated in this study, the goal is to find solutions that “minimize” $\mathbf{f}(C)$ (see Equation 7) in the Pareto-optimality context. Algorithms such as NSGA-II do not return a single “optimal” solution. Instead, return a set of feasible, mutually non-dominated solutions that approximate as much as possible the Pareto set, which represents the optimal trade-off among the multiple objectives [15].
The Decision Maker (DM) must choose the most suitable option from the approximation set to “implement”.
@@ 4.4 Solution Selection
$Q(C)=1-f_{1}-f_{2}$ (12)
This score favors solutions with low intra-community penalty and low inter-community connectivity, as defined in Equation 6, thus promoting partitions that are both cohesive and well-separated. Figure 4 illustrates this behavior: the red dot indicates the solution that maximizes $Q(C)$. We observe that this selection strategy consistently identifies partitions that also perform well under NMI and AMI validation metrics.
@@
Algorithm 3 Solution Selection
1:procedure SelectSolution(Pareto front $\mathcal{F}_{1}$)
2:  Input: Set of non-dominated solutions $\mathcal{F}_{1}=\{p_{1},p_{2},\ldots,p_{k}\}$
3:  Output: Recommended solution $p^{*}$
4:  $\triangleright$ Scalarization-based selection criterion
[…]
6:  $p^{*}\leftarrow\text{null}$
7:  for each solution $p\in\mathcal{F}_{1}$ do
8:   $Q(p)\leftarrow 1-f_{1}(p)-f_{2}(p)$
[…]
11:     $p^{*}\leftarrow p$
12:   end if
13:  end for
14:  $\triangleright$ Domain-specific selection criteria can be applied here
15:  $\triangleright$ e.g., number of key nodes, communities, conductance, etc.
16:  return $p^{*}$
17:end procedure
@@ 4.5 Complexity Analysis
[…]
@@ 4.5.1 Operators Complexity Analysis
[…]
@@
[…]
5 Experimental Results
This section details the methodology and experimental evaluation of HP-MOCD. Our study explores four core dimensions: (i) detection performance on synthetic and real-world networks, (ii) scalability with increasing network size, (iii) parallel efficiency via multithreaded execution, and (iv) robustness under structural noise. Experiments are conducted using standard benchmarks, well-established metrics, and strong baseline algorithms, ensuring fair and reproducible comparisons. To ensure rigorous and interpretable comparisons, we perform pairwise statistical tests (paired t-test with initial $\alpha=0.05$), applying Bonferroni correction to adjust for multiple comparisons. The goal is to determine, for each evaluation, whether HP-MOCD is statistically superior, equivalent, or inferior to the best-performing baseline. We classify each result accordingly and highlight them in bold (superior) in the tables.
5.1 Comparison Algorithms and Parameter Configuration